OpenStudy (anonymous):

consider the line 6x-4y=-6 What is the slope of a line perpendicular to this line? What is the slope of a line parallel to this line?

OpenStudy (anonymous):

perpendicular: y=-2/3x-6

OpenStudy (aroub):

-4y=-6x-6 y=3/2+3/2 parallel= 3/2 perpendicular= -2/3

OpenStudy (anonymous):

parallel: y=3/2x-(any number)

OpenStudy (anonymous):

ax +by = c -> by = c -ax y = c/b -ax/b so slope is -a/b